# Implementation of time-dependent Hartree Fock in real space

### v2

@@ 1 Introduction
Hartree-Fock (HF) theory [1, 2] is a main pillar in quantum chemistry, serving as the starting point for calculating the electronic structure of molecules. HF calculations provide molecular geometry, electron distribution, and ground-state properties [3]. Once the ground state has been calculated, excited states can be calculated using the ground-state results. A first approximation one can make for the excitation energies is the eigenvalue differences of the unoccupied and occupied states, termed the Independent Particle Approximation (IPA) [4]. Configuration Interaction (CI), Equation-of-Motion Coupled Cluster (EOM-CC), and Time-Dependent Hartree-Fock (TDHF) are examples of post-Hartree-Fock methods [5] that can provide more accurate excited states. The exact exchange potential used in HF is orbital-dependent while it is density-dependent in traditional semi-local Kohn-Sham Density Functional Theory (DFT). Hybrid functionals which are a mixture of these two types of exchange potentials are widely used [6, 7] to improve accuracy, so TDHF is a building block for time-dependent density-functional theory (TDDFT) with hybrids [8].
@@
Hartree-Fock theory provides a reasonably accurate description of the ground state of molecular systems, especially for small-to-moderately sized molecules and systems where electron correlation effects are not dominant [8]. Gaussian-type orbitals (GTOs) are typically used as basis sets for Hartree-Fock [9]. Implementation of Hartree-Fock in real space is rare, though a finite-element-type implementation has been published [10]. Using GTOs conveniently allows integrals to be done analytically. However, the nature of orbitals one can obtain in an electronic structure calculations depends on the type and size of the basis set used. The basis set is typically optimized for the ground state which can make description of unoccupied orbitals difficult [11]. This issue is particularly acute for Rydberg states which are highly excited states that are far from the atoms and weakly bound [12], but transitions from valence to Rydberg states can be important in optical properties [13]. To include these states, additional diffuse functions are commonly added to the basis sets [14]. Comparisons of GTOs and real space have shown that such very large, diffuse basis sets can be needed to describe the excited states involved in non-linear optical response [15]. In real space, Rydberg states do not require special diffuse functions but can be captured by systematically converging the physical size of the real-space domain in which the orbitals are calculated. This domain size and the spacing of the grid-points are the parameters controlling the accuracy of a real-space calculation, which can be systematically converged [16]. Other advantages of real-space calculations include favorable scaling for large systems, effective parallelization over domains, and straightforward use of both finite and periodic boundary conditions [16, 17, 18], as well as the ability to handle non-atomic model systems [19]. In real-space however, the calculation of exchange integrals becomes very expensive [20] and these integrals are a major part of any HF calculation underlying TDHF.
@@
In this paper, we present results from our real-space implementation of TDHF in the Octopus code. In section 2, we discuss the general theory behind ground-state HF and TDHF. Section 3.1 discusses the exact exchange potential which is the main feature of HF which sets it apart from traditional Kohn-Sham DFT. As this is the most expensive part of the calculation, we discuss the Adatively Compressed Exchange (ACE) formalism to speed up the exact exchange calculation. We also present the choice of parameters used in our calculations in sections 3.3 and 3.4. In section 4 we discuss the general difficulties in reaching a solution to the HF ground-state in real-space and how to overcome them. Finally, section 4.5 shows the results of TDHF calculations, with comparison to excitation energies obtained from GTOs, demonstrating a good level of agreement and practical attainability of convergence with respect to the real-space grid parameters.
@@ 2 Theory
Ground-state HF calculations involve solving for energy eigenvalues $\epsilon_{i}$ and eigenstates $\{\psi_{i}\}$ with a self-consistent field (SCF) calculation via the Hartree-Fock-Roothan equations [21]:
@@
The four terms in the parentheses are the kinetic energy operator, external potential, Hartree potential, and exchange potential. The Hartree potential depends on the density $\rho\left(r\right)=\sum_{j}^{N_{e}}\left|\psi_{j}\right|^{2}$, where $N_{e}$ is the number of electrons. The exchange potential is used in its exact form and its action on an arbitrary orbital $\psi_{p}$ is given by
@@
Correlation is absent, unlike DFT which has the exchange-correlation functional $V_{\textrm{XC}}$. To calculate the excited-state properties of a system in most linear-response methods, we need extra unoccupied (virtual) orbitals. Self-consistency is not required in generating these unoccupied orbitals because they do not contribute to the density or exchange potential of the system.
@@
$\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{-B}&\mathbf{-A}\\ \end{bmatrix}\begin{bmatrix}\mathbf{X}\\ \mathbf{Y}\\ \end{bmatrix}=\omega\begin{bmatrix}\mathbf{1}&\mathbf{0}\\ \mathbf{0}&\mathbf{1}\\ \end{bmatrix}\begin{bmatrix}\mathbf{X}\\ \mathbf{Y}\\ \end{bmatrix}.$ (3)
The matrix elements in the case of a spin-polarized (unrestricted HF, UHF) calculation are given by
@@
$(pq|rs)$ $=\int\frac{d^{3}\textbf{r}^{\prime}d^{3}\textbf{r}}{\lvert\textbf{r}-\textbf{r}^{{}^{\prime}}\rvert}\phi_{p}^{*}(\textbf{r}^{\prime})\phi_{q}(\textbf{r}^{\prime})\phi_{r}^{*}(\textbf{r})\phi_{s}(\textbf{r}).$
For spin-unpolarized (restricted HF, RHF) calculations, the matrix elements are given by [21]
@@
where $\delta_{M}=2$ and 00 for singlets and triplets respectively.
@@
$\vec{d}_{k}=\sum_{ia}\left<i|\vec{r}|a\right>x_{k,ia}$ (7)
like in TDDFT, where the index $ia$ runs over the pairs of occupied-unoccupied orbitals. The isotropically averaged oscillator strength, in the electric dipole approximation, is then calculated as
$f_{k}=\frac{2m_{e}}{3\hbar^{2}}\omega_{k}\left|\vec{d}_{k}\right|^{2}$ (8)
@@
The $\delta$ function is in practice broadened for plotting with a Lorentzian or Gaussian function.
@@ 3.1 Exchange Potential and Adaptively Compressed Exchange (ACE)
Equation 2 can be solved as a Poisson equation with $\rho(r)=\psi_{j}(r)\psi_{i}(r)$, using various optimized Poisson solvers such as interpolating scaling functions, like for the Hartree potential [29]. The cost of solving equation 2 by direct integration for each $r$ is $\mathcal{O}(N_{\mu}^{2})$, while the cost of solving as a Poisson problem is $\mathcal{O}(N_{\mu}\log N_{\mu})$, where $N_{\mu}$ is the number of grid points. This is a significant improvement, but exact exchange still remains considerably more computationally intensive than a semi-local functional.
@@
$\zeta_{k}(r)=\sum_{i}^{N_{e}}W_{i}(r)(L^{-{\rm T}})_{ik}.$ (11)
Finally, we obtain the ACE operator as
$\hat{V}_{\textrm{X}}^{\textrm{ACE}}(r,r^{\prime})=-\sum_{k=1}^{N_{e}}\zeta_{k}(r)\zeta_{k}(r^{\prime}).$ (12)
@@
ACE significantly reduces the cost of application of the exchange operator by reducing the prefactor but keeps the scaling with grid points the same as the scaling for LDA or HF without ACE [30]. Instead of calculating $\hat{V}_{\textrm{X}}$ by integration from equation 2 for each iteration of the SCF cycle, ACE does it only once and improves it iteratively until the SCF density converges. After the self-consistency is reached, the action of $\hat{V}_{\textrm{X}}^{\textrm{ACE}}$ on an orbital is the same as the action of $\hat{V}_{\textrm{X}}$. This significantly reduces the overall cost of computation because the cost of application of the exchange operator can often be more than $95\%$ of the total cost in HF calculations.
@@
Once the solution to the ground-state is achieved, we calculate a set of unoccupied orbitals via a non-self-consistent calculation. ACE is used here too, with the only difference from the ground-state calculation being the inclusion of unoccupied orbitals in equations 10, 11 and 12. Although one can obtain these unoccupied orbitals in the ground-state SCF by calculating an extra set of eigenfunctions with zero occupation which do not contribute to the total energy, it is preferable to calculate the ground-state without unoccupied orbitals first, since calculation of these extra eigenfunctions in all iterations before convergence is not useful. The number of unoccupied orbitals one can obtain also depends on the size of the basis set: the number of eigenfunctions obtainable is equal to the size of the Hamiltonian matrix. Hence, the number of unoccupied orbitals obtainable from GTOs will be lower than that obtainable from the (orders of magnitude) larger number of real-space grid points. Usually, one calculates all possible unoccupied orbitals with GTOs, but in real space only a small fraction of the possible unoccupied orbitals are calculated.
@@
ACE calculates the ground-state exchange potential iteratively as a whole by reducing the number times the exchange matrix is calculated in the SCF cycle. It is not implemented yet in Octopus to calculate individual exchange integrals with ACE, as would be needed for its use in TDHF. Refs. [33, 34] have shown this is possible with plane-waves and Gaussian basis sets using discrete variable representation [35] schemes. Whether this provides any speedup in real-space is yet to be studied and is a subject for future work. SCF, without ACE, requires solving $N_{e}^{2}$ integrals for each iteration. However, TDA-TDHF requires solving $N_{e}N_{\rm unocc}$ integrals only once, to create the TDHF matrix which is diagonalized. This means the prefactor for TDHF integrals is generally lower than that for the SCF problem, and so it is more necessary to reduce the cost of exchange in SCF than in TDHF.
@@ 3.2 Mixing in SCF cycle
The nonlocality of the exchange potential makes different and more difficult SCF convergence behavior than in DFT. To facilitate faster convergence, different quantities can be mixed together from consecutive iterations and be used in the next iteration. The quantity that is usually mixed is the density, but one can also mix the potential or the wavefunctions. It should also be noted that mixing non-local potentials is not as straightforward as mixing local potentials. In our calculations with potential mixing, we mix only the local part of the potential, i.e. the Hartree potential $V_{\rm H}$.
@@
$q_{N+1}(r)=\alpha q_{N-1}(r)+(1-\alpha)q_{N}(r)$ (13)
@@ 3.3 Pseudopotentials
In real-space or plane-wave calculations, pseudopotentials (PPs) are the standard approach to account only implicitly for the core electrons of a system, which can be done because the core electrons usually do not take part in bonding or electronic and optical properties. In real-space, PPs also solve a crucial problem. The $-Z/r$ Coulomb potential is divergent at $r=0$ and makes the wavefunctions vary very rapidly near the nucleus, requiring an infeasible large number of grid points to describe. PPs circumvent this problem [16]. By contrast, in Gaussian basis sets, usually all the electrons are treated explicitly, except for heavy elements in which case effective core potentials (a form of PPs) [39] are used. PPs are generated from all-electron calculations for a given DFT functional. Typically only local or semi-local functionals are used. Approaches to generate PPs for HF have been developed [40, 41] but such PPs are rarely used in practice [42] and are not supported by Octopus. In this paper, we use optimized norm-conserving Vanderbilt [43] scalar relativistic pseudopotentials from pseudo-dojo [44]. We use LDA [26, 27] pseudopotentials for all HF ground-state calculations, as well as the LDA calculations reported later. For DFT calculations with BLYP [45, 46] for our TDDFT comparison, we have used PBE PPs as BLYP and PBE are both GGA functionals and pseudo-dojo currently doesn’t have BLYP PPs. For these TDDFT calculations, we note that Octopus currently only supports LDA kernels and so the LDA kernel [26, 27] is used for our TDDFT calculations based on a BLYP ground state, as is often done in Octopus calculations [15]. Using PPs generated with one functional for a DFT calculation with a different functional is a common practice in DFT, and ubiquitous for hybrid functionals. A recent study shows that the inconsistent use of PPs does not make very significant differences [47] in DFT.

### v3

# Implementation of time-dependent Hartree-Fock in real space

@@ 1 Introduction
Hartree-Fock (HF) theory [1, 2] is a main pillar in quantum chemistry, serving as the starting point for calculating the electronic structure of molecules. HF calculations provide molecular geometry, electron distribution, and ground-state properties [3]. Once the ground state has been calculated, excited states can be calculated using the ground-state results. A first approximation one can make for the excitation energies is the eigenvalue differences of the unoccupied and occupied orbitals, termed the Independent Particle Approximation (IPA) [4]. Configuration Interaction (CI), Equation-of-Motion Coupled Cluster (EOM-CC), and Time-Dependent Hartree-Fock (TDHF) are examples of post-Hartree-Fock methods [5] that can provide more accurate excited states. The exact-exchange potential used in HF is orbital-dependent while it is density-dependent in traditional semi-local Kohn-Sham Density Functional Theory (DFT). Hybrid functionals which are a mixture of these two types of exchange potentials are widely used [6, 7] to improve accuracy, so TDHF is a building block for time-dependent density-functional theory (TDDFT) with hybrids [8].
@@
Hartree-Fock theory provides a reasonably accurate description of the ground state of molecular systems, especially for small-to-moderately sized closed-shell molecules and systems where electron correlation effects are not dominant [8]. Gaussian-type orbitals (GTOs) are typically used as basis sets for Hartree-Fock [9]. Implementation of Hartree-Fock in real space is rare, though implementations on special grids for atoms and diatomic molecules [10] and a finite-element-type implementation have been published [11]. Using GTOs conveniently allows integrals to be done analytically. However, the nature of orbitals one can obtain in electronic structure calculations depends on the type and size of the basis set used. The basis set is typically optimized for the ground state which can make description of unoccupied orbitals difficult [12]. This issue is particularly acute for Rydberg states which are highly excited states that are far from the atoms and weakly bound [13], but transitions from valence to Rydberg states can be important in optical properties [14]. To include these states, additional diffuse functions are commonly added to the basis sets [15]. Comparisons of GTOs and real space have shown that such very large, diffuse basis sets can be needed to describe the excited states involved in non-linear optical response [16]. In real space, Rydberg states do not require special diffuse functions but can be captured by systematically converging the physical size of the real-space domain in which the orbitals are calculated. This domain size and the spacing of the grid-points are the parameters controlling the accuracy of a real-space calculation, which can be systematically converged [17]. Other advantages of real-space calculations include favorable scaling for large systems, effective parallelization over domains, and straightforward use of both finite and periodic boundary conditions [17, 18, 19], as well as the ability to handle non-atomic model systems [20]. In real-space however, the calculation of exchange integrals becomes very expensive [21] and these integrals are a major part of any HF calculation underlying TDHF.
@@
In this paper, we present results from our real-space implementation of TDHF in the Octopus code. In section 2, we discuss the general theory behind ground-state HF and TDHF. Section 3.1 discusses the exact-exchange potential which is the main feature of HF that sets it apart from Kohn-Sham DFT. As this is the most expensive part of the calculation, we discuss the Adatively Compressed Exchange (ACE) formalism to speed up the exact-exchange calculation. We also present the choice of parameters used in our calculations in sections 3.3 and 3.4. In section 4 we discuss the general difficulties in reaching a solution to the HF ground state in real space and how to overcome them. Finally, section 4.5 shows the results of TDHF calculations, with comparison to excitation energies obtained from GTOs, demonstrating a good level of agreement and practical attainability of convergence with respect to the real-space grid parameters.
@@ 2 Theory
Ground-state HF calculations involve solving for single-particle energy eigenvalues $\{\epsilon_{i}\}$ and corresponding eigenstates $\{\psi_{i}\}$ with a self-consistent field (SCF) calculation via the Hartree-Fock-Roothan equations [22]:
@@
The four terms in the parentheses are the kinetic energy operator, external potential, Hartree potential, and exchange potential. The Hartree potential depends on the density $\rho\left(\textbf{r}\right)=\sum_{j}^{N_{e}}\left|\psi_{j}\left(\textbf{r}\right)\right|^{2}$, where $N_{e}$ is the number of electrons. The exchange potential is used in its exact form and its action on an arbitrary orbital $\psi_{p}$ is given by
@@
Correlation is absent, unlike DFT which has the exchange-correlation potential $V_{\textrm{XC}}$. To calculate the excited-state properties of a system in most linear-response methods, we need extra unoccupied (virtual) orbitals. Self-consistency is not required in generating these unoccupied orbitals because they do not contribute to the density or exchange potential of the system.
@@
$\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{B^{*}}&\mathbf{A^{*}}\\ \end{bmatrix}\begin{bmatrix}\mathbf{X}\\ \mathbf{Y}\\ \end{bmatrix}=\omega\begin{bmatrix}\mathbf{1}&\mathbf{0}\\ \mathbf{0}&-\mathbf{1}\\ \end{bmatrix}\begin{bmatrix}\mathbf{X}\\ \mathbf{Y}\\ \end{bmatrix}.$ (3)
The matrix elements in the case of spin-polarized calculations, based on unrestricted HF (UHF) calculations, are given by
@@
$(pq|rs)$ $=\int\frac{d^{3}\textbf{r}^{\prime}d^{3}\textbf{r}}{\lvert\textbf{r}-\textbf{r}^{{}^{\prime}}\rvert}\phi_{p}^{*}(\textbf{r}^{\prime})\phi_{q}(\textbf{r}^{\prime})\phi_{r}^{*}(\textbf{r})\phi_{s}(\textbf{r}).$ (6)
For spin-unpolarized calculations, based on restricted HF (RHF) ground-state calculations, the matrix elements are given by [22]
@@
where $\delta_{M}=2$ and 00 for singlets and triplets respectively. The exchange integral is $\left(ai|jb\right)$ and the Coulomb integral is $\left(ab|ji\right)$. The factor $\delta_{M}$ comes from the cancellation or addition of the exchange term in triplets $\frac{1}{\sqrt{2}}\left(|\uparrow\downarrow>+|\downarrow\uparrow>\right)$ or singlets $\frac{1}{\sqrt{2}}\left(|\uparrow\downarrow>-|\downarrow\uparrow>\right)$, respectively [24, 25], which also occurs in TDDFT [22, 19] and the Bethe-Salpeter equation [26].
@@
$\textbf{d}_{k}=\sum_{ia}\left<i|\hat{\textbf{r}}|a\right>x_{k,ia}$ (8)
like in TDDFT, where the indices $ia$ run over the pairs of occupied and unoccupied orbitals. The isotropically averaged oscillator strength, in the electric dipole approximation, is then calculated as
$f_{k}=\frac{2m_{e}}{3\hbar^{2}}\omega_{k}\left|\textbf{d}_{k}\right|^{2}$ (9)
@@
The $\delta$-function is in practice broadened for plotting with a Lorentzian or Gaussian function.
@@ 3.1 Exchange Potential and Adaptively Compressed Exchange (ACE)
Equation 2 can be solved as a Poisson equation with $\rho(\textbf{r})=\psi_{j}(\textbf{r})\psi_{i}(\textbf{r})$, using various optimized Poisson solvers such as interpolating scaling functions, like for the Hartree potential. The efficiency of these solvers, including their parallelization, has been studied in detail in the context of Octopus in Ref. [33]. The cost of solving equation 2 by direct integration for each $r$ is $\mathcal{O}(N_{\mu}^{2})$, while the cost of solving as a Poisson problem is $\mathcal{O}(N_{\mu}\log N_{\mu})$, where $N_{\mu}$ is the number of grid points. This is a significant improvement, but exact exchange still remains considerably more computationally intensive than a semi-local functional.
@@
$\zeta_{k}(\textbf{r})=\sum_{i}^{N_{e}}W_{i}(\textbf{r})(L^{-{\rm T}})_{ik},$ (12)
where $-{\rm T}$ indicates the inverse of the transpose. Finally, we obtain the ACE operator as
$\hat{V}_{\textrm{X}}^{\textrm{ACE}}(\textbf{r},\textbf{r}^{\prime})=-\sum_{k=1}^{N_{e}}\zeta_{k}(\textbf{r})\zeta_{k}(\textbf{r}^{\prime}).$ (13)
@@
ACE significantly reduces the cost of application of the exchange operator by reducing the prefactor but keeps the scaling with grid points the same as the scaling for LDA or HF without ACE [34]. Instead of calculating $\hat{V}_{\textrm{X}}$ by integration from equation 2 for each iteration of the SCF cycle, ACE does it only once and improves it iteratively until the SCF density converges. Also, it is important to point out that ACE is not an approximation because after the self-consistency is reached, the action of $\hat{V}_{\textrm{X}}^{\textrm{ACE}}$ on an orbital is the same as the action of $\hat{V}_{\textrm{X}}$. This significantly reduces the overall cost of computation because the cost of application of the exchange operator can often be more than $95\%$ of the total cost in HF calculations.
@@
Once the solution to the ground state is achieved, we calculate a set of unoccupied orbitals via a non-self-consistent calculation. The number of unoccupied orbitals one can obtain depends on the size of the basis set: the number of eigenfunctions obtainable is equal to the size of the Hamiltonian matrix. Hence, the number of unoccupied orbitals obtainable from GTOs will be lower than that obtainable from the (orders of magnitude) larger number of real-space grid points. Usually, one calculates all possible orbitals with GTOs, but in real space only the occupied orbitals (contributing to the Hamiltonian) are calculated in SCF. To generate unoccupied orbitals, a separate non-self-consistent run is performed with a fixed Hamiltonian, calculating only a small fraction of the possible unoccupied orbitals. This is the main difference compared to GTO-based implementations where all possible unoccupied orbitals are usually calculated. Although one can obtain these unoccupied orbitals in the ground-state SCF by calculating an extra set of eigenfunctions with zero occupation which do not contribute to the total energy, it is preferable to calculate the ground state without unoccupied orbitals first, since calculation of these extra eigenfunctions in all iterations before convergence is time-consuming and not useful. Once the number of desired unoccupied orbitals is chosen, the same eigensolver algorithms used for occupied eigenstates in SCF are used to calculate the unoccupied orbitals. Different iterative eigensolvers are available in Octopus such as conjugate gradients [37] and preconditioned Lanczos [38]. In this work, we have used conjugate gradients except in some cases where the highest few unoccupied orbitals would not converge. In this situation, continuing with the preconditioned Lanczos algorithm solved the problem. ACE is used here too to calculate the Hamiltonian matrix, with the only difference from the ground-state calculation being the inclusion of unoccupied orbitals in equations 11, 12 and 13.
@@
ACE calculates the ground-state exchange potential as a whole iteratively by reducing the number of times the exchange matrix is calculated in the SCF cycle. It is not implemented yet in Octopus to calculate individual exchange integrals with ACE, as would be needed for its use in TDHF. Refs. [39, 40] have shown this is possible with plane-waves and Gaussian basis sets using discrete variable representation [41] schemes. Whether this provides any speedup in real space is yet to be studied and is a subject for future work. SCF, without ACE, requires solving $N_{e}^{2}$ integrals for each iteration. However, TDA-TDHF requires solving $N_{e}N_{\rm unocc}$ integrals only once, to create the TDHF matrix which is diagonalized. This means the prefactor for TDHF integrals is generally lower than that for the SCF problem, and so it is more necessary to reduce the cost of exchange in SCF than in TDHF.
@@ 3.2 Mixing in SCF cycle
The nonlocality of the exchange potential leads to different and more difficult SCF convergence behavior than in semi-local DFT. To facilitate faster convergence, different quantities can be mixed together from consecutive iterations and be used in the next iteration [32]. The quantity that is usually mixed is the density, but one can also mix the potential or the wavefunctions. It should also be noted that mixing non-local potentials is not as straightforward as mixing local potentials. In our calculations with potential mixing, we mix only the local part of the potential, i.e. the Hartree potential $V_{\rm H}$.
@@
$q_{N+1}(\textbf{r})=\alpha q_{N-1}(\textbf{r})+(1-\alpha)q_{N}(\textbf{r})$ (14)
@@ 3.3 Pseudopotentials
In real-space or plane-wave calculations, pseudopotentials (PPs) are the standard approach to account only implicitly for the core electrons of a system, which can be done because the core electrons usually do not take part in bonding or electronic and optical properties. In real space, PPs also solve a crucial problem. The $-Z/r$ Coulomb potential is divergent at $r=0$ and makes the wavefunctions vary very rapidly near the nucleus, requiring an infeasibly large number of grid points to describe them. PPs circumvent this problem [17]. By contrast, in Gaussian basis sets, usually all the electrons are treated explicitly, except for heavy elements in which case effective core potentials (a form of PPs) [45] are used. PPs are generated from all-electron calculations for a given DFT functional. Typically only local or semi-local functionals are used. Approaches to generate PPs for HF have been developed [46, 47] but such PPs are rarely used in practice [48] and are not supported by Octopus. In this paper, we use optimized norm-conserving Vanderbilt [49] scalar relativistic pseudopotentials from pseudo-dojo [50]. We use LDA [30, 31] pseudopotentials for all HF ground-state calculations, as well as the LDA calculations reported later. For DFT calculations with BLYP [51, 52] for our TDDFT comparison, we have used PBE PPs as BLYP and PBE are both GGA functionals and pseudo-dojo currently doesn’t have BLYP PPs. For these TDDFT calculations, we note that Octopus currently only supports LDA kernels and so the LDA kernel [30, 31] is used for our TDDFT calculations based on a BLYP ground state, as is often done in Octopus calculations [16]. Using PPs generated with one functional for a DFT calculation with a different functional is a common practice in DFT, and ubiquitous for hybrid functionals. A recent study shows that the inconsistent use of PPs does not make very significant differences [53] in DFT.